# Multi-Hypothesis Prediction for Portfolio Optimization: A Structured Ensemble Learning Approach to Risk Diversification

### v4

@@ Abstract
A framework for portfolio allocation based on multiple hypotheses prediction using structured ensemble models is presented. Portfolio optimization is formulated as an ensemble learning problem, where each predictor focuses on a specific asset or hypothesis. The portfolio weights are determined by optimizing the ensemble’s parameters, using an equal-weighted portfolio as the target, serving as a canonical basis for the hypotheses. Diversity in learning among predictors is parametrically controlled, and their predictions form a structured input for the ensemble optimization model. The proposed methodology establishes a link between this source of learning diversity and portfolio risk diversification, enabling parametric control of portfolio diversification prior to the decision-making process. Moreover, the methodology demonstrates that the diversity in asset or hypothesis selection, based on predictions of future returns, before and independently of the ensemble learning stage, also contributes to the out-of-sample portfolio diversification. The sets of assets with more diverse but lower average return predictions are preferred over less diverse selections. The methodology enables parametric control of diversity in both the asset selection and learning stages, providing users with significant control over out-of-sample portfolio diversification prior to decision-making. Experiments validate the hypotheses across one-step and multi-step decisions for all parameter configurations and the structured model variants using equity portfolios.
@@ 1 Introduction
While the aforementioned solutions reduce parameter uncertainty and can improve out-of-sample performance, they fail to establish a connection between in-sample and out-of-sample diversification, which is a key feature addressed in this work. In this context, a framework is proposed for portfolio optimization using structured ensembles for prediction with multiple hypotheses. Each portfolio component corresponds to a hypothesis, with its return series modeled by an individual predictor. This approach differs from return forecasting (Ma et al., 2021) or scoring models (Nguyen and Lo, 2012; Zhao et al., 2022), where models are used to score assets and predict their returns before applying optimization techniques for portfolio allocation. Building on prior research, the degree of diversity in ensemble learning is parametrically controlled during predictor training and encoded in a structured data set of predictions (Dominguez et al., 2025). This data set is used as input for an ensemble model, with the equal-weight portfolio serving as the target, thereby creating a canonical hypothesis space for portfolio predictions. Scaled ensemble parameters subsequently determine portfolio weights in the allocation problem. To the best of the authors’ knowledge, this is the first plug-in method enabling parametrically controlled portfolio risk diversification prior to the optimization step.
@@
    Introduced a consistent framework for using structured ensemble models in multiple hypothesis prediction settings for portfolio optimization, facilitated by the equal-weighted portfolio target selection.
@@
    Presented a plug-in method with parametric a priori control over the portfolio’s out-of-sample risk diversification level.
@@
    Proposed a strategy to parametrically control the diversity linked to asset or hypothesis selection with out-of-sample portfolio diversification.
@@
    Systematically validated the link between diversity in structured ensembles for multiple hypothesis prediction and risk diversification across all experiments and models, for both one-step and multi-step ahead decision cases.
The paper is organized as follows. Section 2 provides a review of previous work on data-driven and robust optimization, Predict-and-Optimize frameworks, structured ensemble learning, and diversity. Section 3 describes the proposed framework and methodology. Section 4 presents the experimental results and discussions. Finally, Section 5 offers the concluding remarks and future outlook.
@@ 2.1 Data-Driven and Robust Optimization
Optimization under uncertainty is critical in many applications, and two prominent approaches are data-driven optimization and robust optimization. Data-driven optimization relies on data to estimate uncertainties, using statistical models or machine learning to incorporate these estimates into decision-making (Bertsimas et al., 2013; Shapiro et al., 2009). It aims to optimize expected performance metrics such as cost or profit by leveraging historical or real-time data. On the other hand, robust optimization focuses on worst-case scenarios, ensuring decisions remain feasible and effective for all outcomes within a predefined uncertainty set, such as polyhedral or ellipsoidal bounds (Ben-Tal et al., 2009). This approach prioritizes stability and reliability, often at the expense of conservatism if the uncertainty set is poorly calibrated (Bertsimas et al., 2011). Recent advances combine the strengths of both approaches, such as distributionally robust optimization (DRO), which optimizes for the worst-case distribution within plausible distributions, bridging probabilistic and robust methodologies (Esfahani and Kuhn, 2015; Rahimian and Mehrotra, 2022). In the context of portfolio optimization, data-driven approaches use historical and real-time market data to model uncertainties, applying methods such as stochastic programming, machine learning, or empirical scenario generation to predict asset returns and covariance structures (Roncalli, 2013; Giglio et al., 2022). Building on these, this work presents a mixed method that combines both approaches, featuring a plug-in facet but robust in that the risk diversification level can be adjusted parametrically before optimization, thereby addressing the disconnection between out-of-sample diversification and input data in the optimization process.
@@ 3.1 Preliminary
If the aforementioned portfolio optimization can be formulated as a prediction task, then the objective utility function of the mean variance in (1) can be represented as an MSE loss function (Cai et al., 2024). This setup allows for finding the optimal weights that minimize the total prediction error of the portfolio, accounting for both bias and variance, while ensuring that the expected return remains above the target value. In order to incorporate the diversity into such an optimization, this work applies ensemble models to harness diversity in ensemble learning and linking it to portfolio diversification. In a multiple hypothesis prediction setup in which each hypothesis is focused on a particular portfolio constituent, the portfolio can be model as an ensemble of hypotheses so that the diversity term in the Bias-Variance-Diversity decomposition with the MSE Loss can be connected to diversification and controlled parametrically, while the centroid combiner rule is respected in the portfolio target selection. To elaborate, in a supervised setting with the data set of time series returns ($\boldsymbol{x}$, $\boldsymbol{r}$), the Bias-variance-diversity decomposition can be expressed in terms of the MSE loss $\mathcal{L}(\boldsymbol{r},\boldsymbol{\bar{r}})$ as (Wood et al., 2024):
@@
In the next section, the portfolio-structured ensemble model (PSEM) is introduced, where each individual hypothesis predicts a specific portfolio constituent, and the diversity in (2) is linked to portfolio generalization performance and risk diversification. The ensemble targets an equal-weighted portfolio, forming a canonical basis of hypotheses or constituent returns, ensuring that the centroid combiner rule for MSE loss is satisfied.
@@ 3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
The PSEM framework is presented as a tool to model a portfolio as a structured ensemble for prediction across multiple hypotheses. This enables a model-centric approach that links predictive generalization within these models to the out-of-sample risk management of a portfolio. Formally, the PSEM can be defined as a map $\mathcal{E}:\boldsymbol{x}\in\mathbb{R}^{N\times M}\to\boldsymbol{r}\in\mathbb{R}^{N}$ which aggregates M predictions from multiple individual predictors or hypotheses {$f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$, parameterized by $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, to produce portfolio predictions for a given asset selection time series return $\boldsymbol{x}$:
@@
$\delta(f_{\boldsymbol{\theta}_{j}}(x_{ij}))=\begin{cases}1-\varepsilon&\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(x_{ij}),y_{ij})>\mathcal{L}(f_{\boldsymbol{\theta}_{k}}(x_{ik}),y_{ik})\ \ \forall k\neq j\\[10.0pt] \frac{\varepsilon}{M-1}&\text{otherwise}\end{cases}$ (5)
$\delta(f_{\boldsymbol{\theta}_{j}}(x_{ij}))$ indicates whether the output of the $j$-th predictor is the top prediction for the $i$-th instance. A smaller $\varepsilon$ places more emphasis on the predictors that perform the best, while a larger $\varepsilon$ encourages updates to less accurate predictors, fostering greater diversity in the ensemble.
@@
Figure 1 shows a representation of the PSEM optimization, illustrated as a shooting gallery where the shooters represent individual predictors. Each shooter receives inputs such as wind conditions, distance, and target position, aiming to hit their respective targets. In each iteration, the trainer penalizes the best shooter by limiting how much their parameters update during gradient descent, while the remaining shooters update equally. This process introduces diversity in learning through $\varepsilon$. Once training is complete, all shots are gathered into a diversified data set, which can be used as input for structured prediction models, like the Structured Radial Basis Function Network (s-RBFN) shown in this figure and trained using least squares (Dominguez et al., 2025).
@@
Figure 1: Structured data set formation, including diversity ($\varepsilon$), and analytical ensemble training of the s-RBFN model, described in Section 3.3. $f_{\boldsymbol{\theta}_{j}}(x_{ij})$ is the $j^{th}$ shooter (predictor) with input data $x_{ij}$ (for example wind, distance, etc.) and $\boldsymbol{\theta}_{j}$ her shooting skills, $\hat{x}_{ij}$ is the $i^{th}$ shoot from $j^{th}$ shooter and $\boldsymbol{\hat{x}}=\boldsymbol{\hat{x}}(\varepsilon)$ is the full set of shoots after training, $\boldsymbol{r}$ is the portfolio target for the training set (equal-weighted portfolio), and $\boldsymbol{w}$ are the s-RBFN model parameters optimized by least-squares.
In the next section, the PSEM optimization using least squares is detailed for the case of the s-RBFN model. Additionally, the equal-weighted portfolio is presented as the target return in the PSEM optimization for all cases, serving as a canonical basis of hypotheses.
@@
3.3 Multi-Hypothesis Portfolio Optimization with Structured Radial Basis Function Network (s-RBFN)

### v5

@@ Abstract
This work proposes a unified framework for portfolio allocation, covering both asset selection and optimization, based on a multiple-hypothesis predict-then-optimize approach. The portfolio is modeled as a structured ensemble, where each predictor corresponds to a specific asset or hypothesis. Structured ensembles formally link predictors’ diversity — captured via ensemble loss decomposition — to out-of-sample risk diversification. A structured dataset of predictor outputs is constructed with a parametric diversity control, influencing both the training process and diversification outcomes. This dataset feeds a supervised ensemble model whose target portfolio must align with the ensemble combiner rule implied by the loss. For squared loss, the arithmetic mean applies, yielding the equal-weighted portfolio as the optimal target. For asset selection, a novel method is introduced that prioritizes assets from more diverse predictor sets, even at the cost of lower average predicted returns, via a diversity–quality trade-off. This diversity is applied prior to optimization and can integrate into other allocation techniques. Experiments on SP500 stocks validate the theoretical framework and show that both sources of diversity expand the limits of portfolio diversification, achieving strong performance across one-step and multi-step allocation tasks.
@@ 1 Introduction
While the aforementioned solutions reduce parameter uncertainty and can improve out-of-sample performance, they fail to establish a connection between in-sample and out-of-sample diversification, which is a key feature addressed in this work. In this context, a framework is proposed for portfolio optimization using structured ensembles for prediction with multiple hypotheses. Each portfolio component corresponds to a hypothesis, with its return series modeled by an individual predictor. This approach differs from return forecasting (Ma et al., 2021) or scoring models (Nguyen and Lo, 2012; Zhao et al., 2022), where models are used to score assets and predict their returns before applying optimization techniques for portfolio allocation. Building on prior research, the degree of diversity in ensemble learning is parametrically controlled during predictor training and encoded in a structured data set of predictions (Dominguez et al., 2025). This dataset serves as input for an ensemble model designed to predict portfolio returns. The ensemble model is optimized as a supervised prediction problem, using as a target a portfolio whose weights are chosen to match the ensemble combiner rule derived from the Bias-Variance-Diversity decomposition (Wood et al., 2024). In the case of squared loss, the corresponding combiner is the arithmetic mean, implying that the equal-weighted portfolio should be selected. The optimal portfolio weights are then obtained by normalizing the optimal ensemble parameters. To the best of the authors’ knowledge, this is the first plug-in method that enables parametrically controlled portfolio risk diversification prior to the optimization step.
@@
    Introduced a consistent framework for applying structured ensemble models in multi-hypothesis prediction settings for portfolio optimization, where the ensemble combiner rule defines the portfolio target. Specifically, in the case of squared loss, this corresponds to the equal-weighted portfolio.
@@
    Presented a ”plug-in” and robust method enabling parametric a priori control over out-of-sample portfolio risk diversification by linking ensemble diversity, derived from the Bias-Variance-Diversity loss decomposition (Wood et al., 2024), to portfolio risk diversification.
@@
    Proposed a strategy for parametrically controlling diversity in asset or hypothesis selection, leveraging a diversity–quality trade-off in predicted returns to enhance out-of-sample portfolio diversification.
@@
    Systematically validated the expansion of diversification limits (Dalio, 2018) enabled by the proposed framework, as well as through the incorporation of diversity in the asset selection stage across other methodologies, demonstrated empirically in both one-step and multi-step decision settings.
The paper is organized as follows. Section 2 provides a review of previous work on data-driven and robust optimization, Predict-then-Optimize frameworks, structured ensemble learning, and diversity. Section 3 describes the proposed framework and methodology. Section 4 presents the experimental results and discussions. Finally, Section 5 offers the final remarks and future outlook.
@@ 2 Literature Review
Portfolio optimization solutions can be broadly categorized into data-driven approaches and robust optimization methods. Data-driven methods typically rely on plug-in optimization, where the objective function is predefined by frameworks such as Modern Portfolio Theory (MPT), and data is used solely to estimate model parameters. In contrast, robust optimization modifies the objective function and relaxes certain framework assumptions to improve adaptability under uncertainty. This approach was among the first to incorporate loss functions, which were later extended to predictive models and predict-then-optimize frameworks, albeit with computational challenges. Ensemble-based methods further introduce loss functions that account for diversity in predictions. This work integrates ensemble models with multiple hypothesis prediction, establishing a connection between diversity in learning and risk diversification, while also introducing computationally efficient analytical solutions.
@@ 2.1 Data-Driven and Robust Optimization
Optimization under uncertainty is critical in many applications, and two prominent approaches are data-driven optimization and robust optimization. Data-driven optimization relies on data to estimate uncertainties, using statistical models or machine learning to incorporate these estimates into decision-making (Bertsimas et al., 2013; Shapiro et al., 2009). It aims to optimize expected performance metrics such as cost or profit by leveraging historical or real-time data. On the other hand, robust optimization focuses on worst-case scenarios, ensuring decisions remain feasible and effective for all outcomes within a predefined uncertainty set, such as polyhedral or ellipsoidal bounds (Ben-Tal et al., 2009). This approach prioritizes stability and reliability, often at the expense of conservatism if the uncertainty set is poorly calibrated (Bertsimas et al., 2011). Recent advances combine the strengths of both approaches, such as distributionally robust optimization (DRO), which optimizes for the worst-case distribution within plausible distributions, bridging probabilistic and robust methodologies (Esfahani and Kuhn, 2015; Rahimian and Mehrotra, 2022). In the context of portfolio optimization, data-driven approaches leverage historical and real-time market data to model uncertainties, employing methods such as stochastic programming, machine learning, and empirical scenario generation to predict asset returns and covariance structures (Roncalli, 2013; Giglio et al., 2022; Ma et al., 2021; Ozelim et al., 2023). Robust portfolio optimization solutions include fuzzy approaches (Wu and Liu, 2012), genetic algorithms (Akopov, 2014), multiobjective particle swarm optimization (Chen and sha Zhou, 2018), and learning-guided multiobjective evolutionary algorithms (Lwin et al., 2014). Building on these, this work presents a mixed method that combines both approaches, featuring a plug-in facet but robust in that the risk diversification level can be adjusted parametrically before optimization, thus addressing the disconnection between out-of-sample diversification and input data in the optimization process.
@@ 3.1 Preliminary
If the aforementioned portfolio optimization can be formulated as a prediction task, then the objective utility function of the mean-variance in (1) can be represented as an MSE loss function (Cai et al., 2024). This setup allows for finding the optimal weights that minimize the total prediction error of the portfolio, accounting for both bias and variance, while ensuring that the expected return remains above the target value. In order to incorporate the diversity into such an optimization, this work applies ensemble models to harness diversity in ensemble learning and linking it to portfolio diversification. In a multiple-hypothesis prediction setup in which each hypothesis is focused on a particular portfolio constituent, the portfolio can be model as an ensemble of hypotheses so that the diversity term in the Bias-Variance-Diversity decomposition with the MSE Loss can be connected to diversification and controlled parametrically, while the centroid combiner rule is respected in the portfolio target selection. To elaborate, in a supervised setting with the data set of time series returns ($\boldsymbol{x}$, $\boldsymbol{r}$), the Bias-variance-diversity decomposition can be expressed in terms of the MSE loss $\mathcal{L}(\boldsymbol{r},\boldsymbol{\bar{r}})$ as (Wood et al., 2024):
@@
In the next section, the Portfolio-Structured Ensemble Model (PSEM) is introduced, where each individual hypothesis predicts a specific portfolio constituent. The diversity term in equation (2) is linked to both portfolio generalization performance and risk diversification. The predict-then-optimize setting with PSEM for portfolio optimization is then described, in which the target portfolio corresponds to the combiner rule. Under MSE loss, this rule becomes the arithmetic combiner, that is, the equal-weighted portfolio. In this context, the optimal ensemble parameters are equivalent to the optimal portfolio weights.
@@ 3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
The PSEM is a portfolio model formulated as a structured ensemble for prediction across multiple hypotheses. Formally, it can be defined as a map $\mathcal{E}:\boldsymbol{x}\in\mathbb{R}^{N\times M}\to\boldsymbol{r}\in\mathbb{R}^{N}$ which aggregates M predictions from multiple individual predictors or hypotheses {$f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$, parameterized by $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, to produce portfolio predictions for a given asset selection time series return $\boldsymbol{x}$:
@@
$\delta(f_{\boldsymbol{\theta}_{j}}(x_{ij}))=\begin{cases}1-\varepsilon&\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(x_{ij}),y_{ij})<\mathcal{L}(f_{\boldsymbol{\theta}_{k}}(x_{ik}),y_{ik})\ \ \forall k\neq j\\[10.0pt] \frac{\varepsilon}{M-1}&\text{otherwise}\end{cases}$ (5)
$\delta(f_{\boldsymbol{\theta}_{j}}(x_{ij}))$ indicates whether the output of the $j$-th predictor is the top prediction for the $i$-th instance. A smaller $\varepsilon$ places more emphasis on the predictors that perform the best, while a larger $\varepsilon$ encourages updates to less accurate predictors, fostering greater diversity in the ensemble. Algorithm 1 provides the pseudo-code for the structured dataset formation with stochastic gradient descent.
@@
Algorithm 1 Structured Dataset Formation with Gradient Descent
1:Individual predictors $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$ with parameters $\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, input returns $\{x_{i,j}\}^{N,M}$, targets $\{y_{i,j}\}^{N,M}$ with $y_{i,j}=x_{(i+\tau),j}$, learning rates $\eta_{j}$, diversity parameter $\varepsilon$, regularization parameter $\lambda_{p}$, loss function $\mathcal{L}$.
2:Structured dataset $\hat{\boldsymbol{x}}(\varepsilon)$, optimal parameters $\hat{\boldsymbol{\Theta}}$.
3:Randomly initialize $\boldsymbol{\theta}_{j}$ for all $j=1$ to $M$
4:for $i=1$ to $N$ do
5:     for $j=1$ to $M$ do
6:         if $M=1$ then
7:              $\delta(f_{\boldsymbol{\theta}_{j}}(x_{i,j}))\leftarrow 1$
8:         else
9:              for $k=1$ to $M$ do
10:                  if $j\neq k$ then
11:                       if $\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(x_{i,j}),y_{i,j})<\mathcal{L}(f_{\boldsymbol{\theta}_{k}}(x_{i,k}),y_{i,k})$ then
12:                           $\delta(f_{\boldsymbol{\theta}_{j}}(x_{i,j}))\leftarrow 1-\varepsilon$
13:                       else
14:                           $\delta(f_{\boldsymbol{\theta}_{j}}(x_{i,j}))\leftarrow\frac{\varepsilon}{M-1}$
15:                       end if
16:                  end if
17:              end for
18:         end if
19:
$\boldsymbol{\theta}_{j}\leftarrow\boldsymbol{\theta}_{j}-\eta_{j}\left(\frac{\partial\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(x_{i,j}),y_{i,j})}{\partial\boldsymbol{\theta}_{j}}+\frac{\lambda_{p}}{N}\boldsymbol{\theta}_{j}\right)\delta(f_{\boldsymbol{\theta}_{j}}(x_{i,j}))$
20:         $\hat{x}_{i,j}(\varepsilon)\leftarrow f_{\boldsymbol{\theta}_{j}}(x_{i,j})$
21:     end for
22:end for
23:return $\hat{\boldsymbol{x}}(\varepsilon)$ and $\hat{\boldsymbol{\Theta}}$
3.3 Predict-then-optimize Method with PSEM
In the predict-then-optimize setting with PSEM, the optimal portfolio weights $\boldsymbol{w}$ are obtained by minimizing the loss function $\mathcal{L}(\boldsymbol{\hat{r}},\boldsymbol{\bar{r}})$, formulated as a supervised learning problem between the PSEM prediction $\boldsymbol{\hat{r}}$, as defined in Section 3.2, and a portfolio target $\boldsymbol{\bar{r}}$ whose weights reflect the ensemble combiner rule in (2). This combiner rule depends on the choice of loss function, following the bias-variance-diversity decomposition framework proposed by (Wood et al., 2024). For the squared loss, the combiner rule reduces to the arithmetic combiner. When the target portfolio time series is defined using the arithmetic combiner, it is equivalent to the equal-weighted portfolio representing the naive diversification strategy (DeMiguel et al., 2009). The target portfolio returns are given by $\boldsymbol{\bar{r}}=\boldsymbol{x}^{T}\boldsymbol{w}_{\text{target}}=\boldsymbol{x}^{T}\frac{1}{\|\boldsymbol{x}^{T}\|}$, and the optimization is formulated as:
$\min_{\boldsymbol{w}}\left\|\boldsymbol{\bar{r}}-\boldsymbol{\hat{x}}(\varepsilon)^{T}\boldsymbol{w}\right\|^{2}+\lambda_{s}\|\boldsymbol{w}\|_{p}$ (8)
Here, $\lambda_{s}$ denotes the PSEM regularization parameter. The predict-then-optimize approach can be implemented in a ”plug-in” format, where the structured dataset $\boldsymbol{\hat{x}}(\varepsilon)^{T}$ is constructed independently prior to training the structured ensemble and deriving the portfolio weights. In the experimental section, portfolio weight constraints are applied post-optimization, including non-negativity constraints, $w_{i}\geq 0,\ \forall i\in\{1,\ldots,M\}$, and normalization, $\sum_{i=1}^{M}w_{i}=1$. However, the method can also be extended to a fully constrained plug-in optimization framework using quadratic solvers, which is left for future work.
@@ 3.3 Predict-then-optimize Method with PSEM
Figure 1 shows a representation of the PSEM optimization, illustrated as a shooting gallery where the shooters represent individual predictors. Each shooter receives inputs such as wind conditions, distance, and target position, aiming to hit their respective targets. In each iteration, the trainer penalizes the best shooter by limiting how much their parameters update during gradient descent, while the remaining shooters update equally. This process introduces diversity in learning through $\varepsilon$. Once the training of the individual predictors is completed, all shots are gathered into a diversified data set, which can be used as input for structured prediction models, like the Structured Radial Basis Function Network (s-RBFN) (Dominguez et al., 2025), shown in this figure and trained using least squares.
@@
Figure 1: Structured data set formation, including diversity ($\varepsilon$), and analytical ensemble training of the s-RBFN model, described in Section 3.4. $f_{\boldsymbol{\theta}_{j}}(x_{ij})$ is the $j^{th}$ shooter (predictor) with input data $x_{ij}$ (for example wind, distance, etc.) and $\boldsymbol{\theta}_{j}$ her shooting skills, $\hat{x}_{ij}$ is the $i^{th}$ shoot from $j^{th}$ shooter and $\boldsymbol{\hat{x}}=\boldsymbol{\hat{x}}(\varepsilon)$ is the full set of shoots after training, $\boldsymbol{\bar{r}}$ is the portfolio target for the training set (equal-weighted portfolio), and $\boldsymbol{w}$ are the s-RBFN model parameters optimized by least-squares.
In the next section, the PSEM optimization is detailed for the case of the s-RBFN model.
@@
3.4 Example with Structured Radial Basis Function Network (s-RBFN)